OpenStudy (anonymous):

algebra 2 help look below

OpenStudy (anonymous):

\[5\sqrt{x-6}= 20\]

OpenStudy (anonymous):

I'm doing Algebra 1 but lets see...

OpenStudy (anonymous):

I don't think i learned this in algebra 1.. so i dont think u'll know it

OpenStudy (anonymous):

What are these equations called?

OpenStudy (otonogold):

I'm in algebra 2 , Give me a sec lol

OpenStudy (anonymous):

ok

OpenStudy (otonogold):

You have a good calculator for math? If you so you try plugging in numbers for x to see if the outcome is 20 ?

OpenStudy (anonymous):

i dont know how to do that on my scientific calculator

OpenStudy (otonogold):

is it multiple choice?

OpenStudy (anonymous):

anyway, i have to know how to work it out on paper

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

-10, 10, 22, and no solution

OpenStudy (anonymous):

this is just a practice problem so it tells u which it is after

OpenStudy (anonymous):

i did it and got 22, but it said try squaring both sides first..

OpenStudy (anonymous):

then it said the right answer is no solution

OpenStudy (otonogold):

Hold on one sec.

OpenStudy (otonogold):

This is a square root function right?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

this is a radical equation

OpenStudy (anonymous):

thats not how u do it

OpenStudy (anonymous):

its not a function at all.. and i dont need to graph anything

OpenStudy (anonymous):

the steps are: isolate the variable, square both sides, solve for the variable, and check ur work

OpenStudy (otonogold):

Ohhhh I feel slow those are easier lol

OpenStudy (otonogold):

Okay so another example problem say you have √x+7 - 5=x √x+7 = x +5 Isolate the radical (√x+7)^2 = (x+5)^2 Square each other x+7= x^2 + 10x + 25 Simplify 0= x^2+9x+18 Combine like terms 0= (x+3)(x+6) factor x= -3 or x = -6 Zero- Product Property

OpenStudy (anonymous):

thats a different one ok. mine is harderr. because there is a number in front of the radical.

OpenStudy (otonogold):

so that problem the answer was -3? to check your work you can do √ x+7 -5 = x √-3+7-5= -3 √4-5 = -3 2-5= -3 -3=-3

OpenStudy (otonogold):

You still do the same thing if the number is in front but basically you just plug in numbers in your scientific calculator too, that is a lot easier. I know schools or homeschooling allowing you to have that and use it.

OpenStudy (anonymous):

no, technically we don't do that

OpenStudy (anonymous):

we work it out, otherwise, how will u learn?

OpenStudy (anonymous):

5sqrt(x-6) = 20 divide 5 both sides sqrt(x-6) = 4 square both sides x-6 = 16 add 6 both sides x = 22

OpenStudy (anonymous):

there^ that is how u do it

OpenStudy (otonogold):

That is true but we have technology to help

OpenStudy (anonymous):

the practice problems are probably just wrong so.. yea

OpenStudy (anonymous):

otono

OpenStudy (anonymous):

yea, i know but i dont even know how to use it

OpenStudy (anonymous):

cupcake

OpenStudy (anonymous):

i dont know how to do these kind of problems on a calculator

OpenStudy (anonymous):

what

OpenStudy (otonogold):

They didn't teach you ?

OpenStudy (anonymous):

no..

OpenStudy (otonogold):

Dang :(

OpenStudy (anonymous):

doing ur work on a calculator is not encouraged.

OpenStudy (anonymous):

fan me cupcake

OpenStudy (anonymous):

is that all u wanted -.-

OpenStudy (anonymous):

no

OpenStudy (otonogold):

I see but you have to know how to work that calculator, because when you get to trig and pre cal it gets harder and you need that , trust me

OpenStudy (anonymous):

oh, btw, im not gonna take pre cal

OpenStudy (anonymous):

my career is not gonna involve math or science and my mom said i dont need to take it

OpenStudy (otonogold):

Lucky :( lol

OpenStudy (otonogold):

Im sorry I wasn't really help.

OpenStudy (anonymous):

its ok